# VoMP: Predicting Volumetric Mechanical Property Fields

### v1

@@ E.2 Dataset Statistics
Our dataset comprises a diverse collection of 1,692 objects, sourced from four datasets: simready nvidia_omniverse_simready, residential nvidia_omniverse_residential_assets, vegetation nvidia_omniverse_vegetation_assets, and commercial nvidia_omniverse_commercial_assets. As shown in Tb.˜8, the majority of objects belong to the simready and residential categories, with vegetation and commercial objects providing additional variety. Each object is decomposed into multiple segments, with a total of 8,128 segments across the dataset. Most parts are labeled with English material names, and for few parts that do not have these material names, we infer these from the PBR texture names that were applied to these parts. To characterize the physical realism and diversity of the dataset, we analyze the distribution of key material properties for all segments in Tb.˜9. The wide range of material properties highlights the heterogeneity of the dataset, which is essential for robust learning and evaluation of material-aware models. We summarize the most frequent material categories (e.g., metal, plastic, wood, cardboard) and object classes (e.g., residential, shelf, container), along with their respective counts and proportions in Tb.˜10.
We report the statistics of our Material Triplet dataset in Tb.˜11. To train MatVAE, we use the "Filtered Dataset".
@@ F.3 Training
We present our voxelization scheme for training on meshes in Algorithms˜2 and 3. We present the hyperparameters used for training MatVAE and Geometry Transformer in Tb.˜12.
@@ F.5 Baselines
The baselines NeRF2Physics zhai2024physicalpropertyunderstandinglanguageembedded and PUGS shuai2025pugszeroshotphysicalunderstanding in their methods sample points from the NeRF or Gaussian splat, respectively, and predict mechanical properties at those points. To ensure fair comparisons in Tb.˜2 and 6, we explicitly make these methods work on the same set of points in the object on which our method is evaluated.
@@ Appendix G Additional Details on the Simulations
We experiment with Simplicits 10.1145/3658184, a reduced-order simulator (Fig.˜1, 5, 8 and 8) and an accurate finite-element method (FEM) simulator (Fig.˜1, 5 and 8) with our material properties. We also use a FEM simulator for our experiments on interpreting errors in properties (§˜D.4). We use a material point method (MPM) mpm, and an Extended Position Based Dynamics (XPBD) xpbd simulator for our experiments to compare between simulators (Fig.˜2). We share details on these simulations. We also share details on the interpolation we use across all our simulations. We share the hyperparameters used for all the FEM simulations in Tb.˜13.
@@ G.4 Preparing Scenes and Assigning Materials for the Simplicits Solver
G.5 Deforming Splats and Rebndering Deformed Splats

### v2

@@ E.2 Dataset Statistics
Our dataset comprises a diverse collection of 1,692 objects, sourced from four datasets: simready nvidia_omniverse_simready, residential nvidia_omniverse_residential_assets, vegetation nvidia_omniverse_vegetation_assets, and commercial nvidia_omniverse_commercial_assets. As shown in Tb.˜10, the majority of objects belong to the simready and residential categories, with vegetation and commercial objects providing additional variety. Each object is decomposed into multiple segments, with a total of 8,128 segments across the dataset. Most parts are labeled with English material names, and for few parts that do not have these material names, we infer these from the PBR texture names that were applied to these parts. To characterize the physical realism and diversity of the dataset, we analyze the distribution of key material properties for all segments in Tb.˜11. The wide range of material properties highlights the heterogeneity of the dataset, which is essential for robust learning and evaluation of material-aware models. We summarize the most frequent material categories (e.g., metal, plastic, wood, cardboard) and object classes (e.g., residential, shelf, container), along with their respective counts and proportions in Tb.˜12.
We report the statistics of our Material Triplet dataset in Tb.˜13. To train MatVAE, we use the "Filtered Dataset".
@@ F.3 Training
We present our voxelization scheme for training on meshes in Algorithms˜2 and 3. We present the hyperparameters used for training MatVAE and Geometry Transformer in Tb.˜14.
@@ F.5 Baselines
The baselines NeRF2Physics zhai2024physicalpropertyunderstandinglanguageembedded and PUGS shuai2025pugszeroshotphysicalunderstanding in their methods sample points from the NeRF or Gaussian splat, respectively, and predict mechanical properties at those points. To ensure fair comparisons in Tb.˜4 and 6, we explicitly make these methods work on the same set of points in the object on which our method is evaluated.
@@ Appendix G Additional Details on the Simulations
We experiment with Simplicits 10.1145/3658184, a reduced-order simulator (Fig.˜1, 5, 8 and 8) and an accurate finite-element method (FEM) simulator (Fig.˜1, 5 and 8) with our material properties. We also use a FEM simulator for our experiments on interpreting errors in properties (§˜D.4). We use a material point method (MPM) mpm, and an Extended Position Based Dynamics (XPBD) xpbd simulator for our experiments to compare between simulators (Fig.˜2). We share details on these simulations. We also share details on the interpolation we use across all our simulations. We share the hyperparameters used for all the FEM simulations in Tb.˜15.
@@ G.4 Preparing Scenes and Assigning Materials for the Simplicits Solver
G.5 Deforming Splats and Rendering Deformed Splats